OpenStudy (henriquerb):

Someone knows how to integrate sqrt(1+x³)dt?

OpenStudy (kainui):

Since you're integrating with respect to dt, you can pull sqrt(1+x^3) outside the integral and the integral is just: \[\sqrt{(1+x^3)} \int\limits_{}^{}dt=t \sqrt{(1+x^3)}+C\]

OpenStudy (anonymous):

@Kainui : I think he meant dx. Could be a typo.

OpenStudy (kainui):

In that case, good luck. http://www.wolframalpha.com/input/?i=sqrt%281%2Bx%C2%B3%29dx

OpenStudy (anonymous):

Fun stuff bro! >.< .